Donald Trump has written a self improvement book that has the following cost characteristics:

Selling Price $14.00 per book
Variable Cost per unit:
Production $ 6.00
Selling and Administration $ 3.00
Fixed Costs:
Production $100,000 per year
Selling and Administration $20,000 per year

What is the cost formula for total cost for this company, where TC = Total cost and Q = units sold?
Answer #1

Total Cost = Variable Cost + Fixed Cost

Variable Cost = Units sold X Variable cost per unit

Total Variable Cost = (6 + 3) = $ 9 per unit

Fixed Cost = (100,000 + 20,000) = $ 120,000

TC = 9Q + 120,000